# Determination of the stably free cancellation property for orders

### v2

@@ 12.4. A probabilistic algorithm
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Let $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and assume that $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ satisfies $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$. If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma^{\prime})$, then $X$ is free over $\Lambda$.
@@
Suppose that $A_{1}$ satisfies the Eichler condition and that every simple component of $A_{2}$ is a totally definite quaternion algebra. Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. If the following algorithm returns $\mathsf{true}$, then $X$ is free over $\Lambda$; if it returns $\mathsf{inconclusive}$, then nothing can be said about whether $X$ is free over $\Lambda$ or not.
@@
      If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma^{\prime})$, return $\mathsf{true}$.
@@ 13. Further results on SFC for integral group rings
$\tilde{T}\times C_{2},\,\tilde{T}\times Q_{12},\,\tilde{T}\times Q_{20},\,Q_{8}\rtimes\tilde{T},\,Q_{8}\rtimes Q_{12},\,\tilde{I}\times C_{2},\,G_{(192,183)},\,G_{(384,580)},\,G_{(480,962)}.$ (13.1)
@@
For $G=Q_{8}\rtimes Q_{12}$, $\tilde{T}\times Q_{12}$, $\tilde{T}\times Q_{20}$, or $\tilde{I}\times C_{2}$, there exists a normal subgroup $N$ of order $2$, such that $H:=G/N$ is isomorphic to $\mathrm{GL}_{2}(\mathbb{F}_{3})$, $G_{(144,127)}$, $G_{(240,108)}$, or $C_{2}\times A_{5}$, respectively. In each case, $\mathbb{Q}[H]$ satisfies the Eichler condition and this implies that $\mathbb{Q}[H\times C_{2}]$ does as well. Thus $\mathbb{Z}[H\times C_{2}]$ has LFC by Theorem 3.4. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. The cardinalities of the locally free class groups of $\mathbb{Z}[H]$, $\mathbb{Z}[G]$, $\mathbb{Z}[H\times C_{2}]$ and $\Gamma$ can be computed using the the algorithm of [BB06, §3], and thus it is easily verified that equality (6.3) holds. Therefore Proposition 6.10 shows that $\mathbb{Z}[G]$ has SFC if only if $\Gamma$ has SFC.
@@
Therefore, for $G=Q_{8}\rtimes Q_{12}$, $Q_{8}\rtimes\tilde{T}$, $\tilde{T}\times Q_{12}$, $\tilde{T}\times Q_{20}$, or $\tilde{I}\times C_{2}$, we have shown that $\mathbb{Z}[G]$ has SFC if and only if $\Gamma$ has SFC. This is highly advantageous from a computational point of view, since $\dim_{\mathbb{Q}}\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=\frac{1}{2}\dim_{\mathbb{Q}}\mathbb{Q}[G]=\frac{1}{2}|G|$. To verify that $\Gamma$ has SFC, we again apply Algorithm 10.3, this time with the fibre product induced by the decomposition $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=B_{1}\oplus B_{2}$, with $B_{1}$ the largest component of $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ satisfying the Eichler condition (equivalently, with $B_{2}$ the largest component that is the direct sum of totally definite quaternion algebras). Computations in $\overline{\Lambda}^{\times}$ are performed by first computing an explicit embedding into the symmetric group $\mathfrak{S}_{d}$ for some $d\in\mathbb{Z}_{>0}$ (see §11.3). This is less efficient than working with an embedding into $\mathrm{GL}_{n}(\mathbb{F}_{p})$ for appropriate choices of $n$ and $p$, but such an embedding does not exist in general.
@@
$\mathbb{Q},\,\mathbb{Q}(i),\,\mathbb{Q}(\sqrt{-3}),\,\mathbb{Q}(\sqrt{5}),\,\mathbb{Q}(\sqrt{2}),\,\mathbb{Q}(\sqrt{-2}),\,\mathbb{Q}(\zeta_{8}),\,\mathbb{Q}(\sqrt{-15}),\,\mathbb{Q}(i,\sqrt{3}),\,\mathbb{Q}(\sqrt{-3},\sqrt{5}).$
@@
We use the GAP [GAP24] small groups database [BEO02], which in particular, contains all groups $G$ with $|G|\leq 1023$. For each such $G$, we perform the following checks, which can easily be implemented as a program in GAP [GAP24] or Magma [BCP97].
@@
  • If $G\cong H\times K$ where $H$ is any group listed in Theorem 13.1 and $K$ is any of $C_{3},C_{5},C_{7},C_{9}$ or $C_{3}^{2}$, then $\mathbb{Z}[G]$ has SFC by Theorems 13.1 and 4.8.
The groups $G$ with $|G|\leq 1023$ that do not satisfy any of the above conditions are listed in (a)–(g) above. ∎

### v3

@@ 12.4. A probabilistic algorithm
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Let $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and assume that $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ satisfies $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$. If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\bmod S^{\prime}\in\mathrm{im}(\sigma^{\prime})$, then $X$ is free over $\Lambda$.
@@
Suppose that $A_{1}$ satisfies the Eichler condition and that every simple component of $A_{2}$ is a totally definite quaternion algebra. Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. If the following algorithm returns $\mathsf{true}$, then $X$ is free over $\Lambda$; if it returns $\mathsf{inconclusive}$, then nothing can be said about whether $X$ is free over $\Lambda$.
@@
      If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma^{\prime})\bmod S^{\prime}$, return $\mathsf{true}$.
@@ 13. Further results on SFC for integral group rings
(13.1) $\tilde{T}\times C_{2},\,\tilde{T}\times Q_{12},\,\tilde{T}\times Q_{20},\,Q_{8}\rtimes\tilde{T},\,Q_{8}\rtimes Q_{12},\,\tilde{I}\times C_{2},\,G_{(192,183)},\,G_{(384,580)},\,G_{(480,962)}.$
@@
Further details on the computations used in this proof can be found in Appendix A.
@@
For $G=Q_{8}\rtimes Q_{12}$, $\tilde{T}\times Q_{12}$, $\tilde{T}\times Q_{20}$, or $\tilde{I}\times C_{2}$, there exists a normal subgroup $N$ of order $2$, such that $H:=G/N$ is isomorphic to $\mathrm{GL}_{2}(\mathbb{F}_{3})$, $G_{(144,127)}$, $G_{(240,108)}$, or $C_{2}\times A_{5}$, respectively. In each case, $\mathbb{Q}[H]$ satisfies the Eichler condition and this implies that $\mathbb{Q}[H\times C_{2}]$ does as well. Thus $\mathbb{Z}[H\times C_{2}]$ has LFC by Theorem 3.4. Let $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$. The orders of the locally free class groups of $\mathbb{Z}[H]$, $\mathbb{Z}[G]$, $\mathbb{Z}[H\times C_{2}]$ and $\Gamma$ are shown in the table below, and from this it is easily verified that equality (6.2) holds in each case. Therefore Proposition 6.10 shows that $\mathbb{Z}[G]$ has SFC if only if $\Gamma$ has SFC.
[…]
$Q_{8}\rtimes Q_{12}$ $\mathrm{GL}_{2}(\mathbb{F}_{3})$ $^{9} 22$ $2^{7}$ $2^{3}$
$\tilde{T}\times Q_{12}$ $G_{(144,127)}$ $2^{28}\cdot 3^{7}$ $^{5}\cdot 3^{3} 22$$^{23}\cdot 3^{7}$ $2^{10}\cdot 3^{3}$
$\tilde{T}\times Q_{20}$ $G_{(240,108)}$ $2^{41}\cdot 3^{3}$ $2^{9}\cdot 3$ $2^{34}\cdot 3^{3}$ $2^{16}\cdot 3$
$\tilde{I}\times C_{2}$ $C_{2}\times A_{5}$ $^{10} 22$ $2^{6}$ $2^{5}$
@@
Therefore, for $G=Q_{8}\rtimes Q_{12}$, $Q_{8}\rtimes\tilde{T}$, $\tilde{T}\times Q_{12}$, $\tilde{T}\times Q_{20}$, or $\tilde{I}\times C_{2}$, we have shown that $\mathbb{Z}[G]$ has SFC if and only if $\Gamma$ has SFC. This is highly advantageous from a computational point of view, since $\dim_{\mathbb{Q}}\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=\frac{1}{2}\dim_{\mathbb{Q}}\mathbb{Q}[G]=\frac{1}{2}|G|$. To verify that $\Gamma$ has SFC, we again apply Algorithm 10.3, this time with the fiber product induced by the decomposition $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma=B_{1}\oplus B_{2}$, with $B_{1}$ the largest component of $\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ satisfying the Eichler condition (equivalently, with $B_{2}$ the largest component that is the direct sum of totally definite quaternion algebras). Computations in $\overline{\Lambda}^{\times}$ are performed by first computing an explicit embedding into the symmetric group $\mathfrak{S}_{d}$ for some $d\in\mathbb{Z}_{>0}$ (see §11.3). This is less efficient than working with an embedding into $\mathrm{GL}_{n}(\mathbb{F}_{p})$ for appropriate choices of $n$ and $p$, but such an embedding does not exist in general.
@@
$\mathbb{Q},\quad\mathbb{Q}(i),\quad\mathbb{Q}(\sqrt{-3}),\quad\mathbb{Q}(\sqrt{5}),\quad\mathbb{Q}(\sqrt{2}),\quad\mathbb{Q}(\sqrt{-2}),$
$\mathbb{Q}(\zeta_{8}),\quad\mathbb{Q}(\sqrt{-15}),\quad\mathbb{Q}(\zeta_{12}),\quad\mathbb{Q}(\sqrt{-3},\sqrt{5}),\quad\mathbb{Q}(\zeta_{16}).$
@@
We use the GAP [GAP24] small groups database [BEO02], which in particular, contains all groups $G$ with $|G|\leq 1023$. For each such $G$, we have performed the following checks using a computer program whose implementation is discussed in Appendix A.
  • If $m_{\mathbb{H}}(G)\leq 1$ then $\mathbb{Z}[G]$ has SFC by Theorem 4.13.
@@
  • If $G\cong H\times K$ where $H$ is any group listed in (13.1) and $K$ is any of $C_{3},C_{5},C_{7},C_{9}$ or $C_{3}^{2}$, then $\mathbb{Z}[G]$ has SFC by Theorems 13.1 and 4.8.
The program outputs the groups $G$ with $|G|\leq 1023$ for which the question of whether $\mathbb{Z}[G]$ has SFC is not determined by any of the above conditions. These groups are listed in (a)–(g) above. ∎
@@
Appendix A Implementation of code including data and timings
All code used in this article is available at github.com/thofma/TestingSFC.jl/. All timings below are in seconds and are for computations performed using a single core of a server with an Intel Xeon 6226R 2.90GHz CPU and 384GB RAM. The software used was OSCAR 1.4.1 [OSC25, DEF+25] (which includes Hecke [FHHJ17] and GAP [GAP24]) and Magma V2.28-7 [BCP97].
A.1. Computations of locally free class groups
Locally free class groups of orders can be computed using the algorithm of [BB06, §3] implemented in OSCAR. The timings of the computations used in the proofs of Theorems 6.6 and 13.1 are given in Table LABEL:tab:classgroups. Recall that in each case apart from that in the first row, we have $H=G/N$ and $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$, where $N$ is a certain normal subgroup of $G$ of order $2$.
Table 1. Locally free class group computations
$G$ $H$ $\mathrm{Cl}(\mathbb{Z}[G])$ $\mathrm{Cl}(\mathbb{Z}[H])$ $\mathrm{Cl}(\mathbb{Z}[H\times C_{2}])$ $\mathrm{Cl}(\Gamma)$
$\tilde{T}$ 3s 3s
$Q_{8}\rtimes Q_{12}$ $\mathrm{GL}_{2}(\mathbb{F}_{3})$ 17s 1s 19s 1s
$\tilde{T}\times Q_{12}$ $G_{(144,127)}$ 2012s 112s 2205s 115s
$\tilde{T}\times Q_{20}$ $G_{(240,108)}$ 6446s 604s 9337s 1479s
$\tilde{I}\times C_{2}$ $C_{2}\times A_{5}$ 769s 68s 1027s 85s
A.2. The proof of Theorem 9.4
In Table 2 we present details on the computations, which were performed using OSCAR. Recall that for each group $G$ under consideration, we find a quotient $\Gamma$ of $\mathbb{Z}[G]$, for which failure of SFC is shown using Algorithm 9.1. We report the dimension of $\mathbb{Q}\Gamma\cong\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ and the overall run time.
Table 2. Proof of failure of SFC for $\mathbb{Z}[G]$ using Algorithm 9.1
$G$ $\dim(\mathbb{Q}\Gamma)$ total time
$Q_{8}\times C_{2}$ 16 121s
$Q_{12}\times C_{2}$ 16 131s
$Q_{16}\times C_{2}$ 24 136s
$Q_{20}\times C_{2}$ 24 234s
$\tilde{T}\times C_{2}^{2}$ 28 210s
$\tilde{O}\times C_{2}$ 48 5740s
$\tilde{I}\times C_{2}^{2}$ 36 1443s
$G_{(32,14)}$ 20 131s
$G_{(36,7)}$ 20 133s
$G_{(64,14)}$ 24 152s
$G_{(100,7)}$ 24 124s
A.3. The proof of Theorem 13.1
We present details on the computations, which were performed using both OSCAR and Magma. Table 3 contains information on the application of Algorithm 10.3 to the respective integral group ring $\Lambda=\mathbb{Z}[G]$. Recall that in each case, we have $H=G/N$ and $\Gamma=\mathbb{Z}[G]/\mathrm{Tr}_{N}\mathbb{Z}[G]$, where $N$ is a certain normal subgroup of $G$ of order $2$. The column labeled ‘reduced to’ refers to the order, for which the SFC property is shown to be equivalent to the SFC property of $\mathbb{Z}[G]$. The columns grouped together under ‘(9)–(10)’ refer to computations of $\overline{\Lambda}^{\times\mathrm{ab}}$ in Steps (9)–(10) of Algorithm 10.3. (Recall that in the notation of Algorithm 10.3, we have that $\overline{\Lambda}_{2}\cong\overline{\Lambda}$.) We give a description of the finite ring $\overline{\Lambda}$, a list $d_{1}^{n_{1}},\dotsc,d_{r}^{n_{r}}]$ such that $\overline{\Lambda}^{\times\mathrm{ab}}$ is isomorphic to
$\mathbb{Z}/d_{1}^{n_{1}}\mathbb{Z}\times\dotsb\times\mathbb{Z}/d_{r}^{n_{r}}\mathbb{Z},$
and the run time of these steps, respectively. The columns grouped together under ‘(13)–(16)’ refer to freeness checks of test lattices in Steps (13)–(16) of Algorithm 10.3. We report the number $r$ of test lattices checked for freeness and the run time for these steps respectively. Finally, the column labeled ‘total time’ contains the total run time of Algorithm 10.3.
Table 4 contains information on the proof of the SFC property for $\Gamma$ using Algorithm 10.3, where $\Gamma$ is the order for which the SFC property is equivalent to the SFC property of $\mathbb{Z}[G]$. The fiber product used is obtained from the ‘maximal Eichler splitting’ of $\mathbb{Q}\Gamma\cong\mathbb{Q}\otimes_{\mathbb{Z}}\Gamma$ which is induced by the decomposition $\mathbb{Q}\Gamma=B_{1}\oplus B_{2}$, with $B_{1}$ the largest component of $\mathbb{Q}\Gamma$ satisfying the Eichler condition. The columns are similar to the columns of Table 3, except that we now also include information on Step (7), which checks whether the projection $\Lambda_{2}$ of $\Lambda$ onto $B_{2}$ has SFC. We report both $\dim(\mathbb{Q}\Lambda_{2})=\dim(B_{2})$ and the run time of this step.
Table 3. Reduction of the SFC property for $\Lambda=\mathbb{Z}[G]$ via Algorithm 10.3
(9)–(10) (13)–(16)
$G$ $H$ reduced to $\overline{\Lambda}$ $\overline{\Lambda}^{\times\mathrm{ab}}$ time $r$ time total time
$Q_{8}\rtimes\tilde{T}$ $G_{(96,204)}$ $\Gamma$ $\mathbb{F}_{2}[H]$ $2^{6},12]$ 625s 2 362s 1080s
$G_{(192,183)}$ $Q_{8}\rtimes Q_{12}$ $\mathbb{Z}[H]$ $\mathbb{F}_{2}[H]$ $2^{10},4,8]$ 686s 4 810s 1612s
$G_{(480,962)}$ $Q_{8}\rtimes Q_{12}$ $\mathbb{Z}[H]$ $\mathbb{F}_{2}[H]$ $4^{12},24^{2}]$ 636s 11 67730s 104619s
$G_{(384,580)}$ $G_{(192,183)}$ $\mathbb{Z}[H]$ $\mathbb{F}_{2}[H]$ $2^{17},4^{4},8^{2}]$ 38750s 10 107968s 152376s
Table 4. Proof of SFC property for $\Lambda=\Gamma$ using Algorithm 10.3 and the ‘maximal Eichler splitting’
(9)–(10) (13)–(16) (7)
$G$ $\dim(\mathbb{Q}\Gamma)$ $\overline{\Lambda}^{\times\mathrm{ab}}$ time $r$ time $\dim(\mathbb{Q}\Lambda_{2})$ time total time
$Q_{8}\rtimes Q_{12}$ 48 $2^{2},4,8^{2}]$ 51s 5 9s 12 27s 201s
$\tilde{T}\times Q_{12}$ 144 $2^{6},6^{3},24]$ 16s 4 380s 12 27s 532s
$\tilde{T}\times Q_{20}$ 240 $2^{8},4^{4},12,14]$ 10183s 8 6800s 12 255s 19098s
$\tilde{I}\times C_{2}$ 120 $2^{2},4^{2},8,24]$ 28s 5 210s 16 88s 485s
$Q_{8}\rtimes\tilde{T}$ 96 $2^{5},12]$ 5100s 2 480s 8 9s 7055s
A.4. The proof of Theorem 13.4
There are two implementations of the checks described in this proof. One is written in Magma and the other written in OSCAR. For technical reasons, there are in fact two separate programs in each case: one for $|G|=512$, and another for $|G|\leq 1023$ and $|G|\neq 512$.